root node left root child nodeA child 1 nodeA child 2 nodeA child 3 right root child nodeB child 1

root node left root child nodeA child 1 nodeA child 2 nodeA child 3 right root child nodeB child 1

root node left root child nodeA child 1 nodeA child 2 nodeA child 3 right root child nodeB child 1